if your in the bed of a truck doing 50 mph and you throw a baseball from the Back of the truck and you throw a 50 mph pitch is the Baseball traveling at 100 mph?
1
Expert's answer
2019-06-20T11:38:13-0400

1)if the baseball is flying in the direction of the truck, baseball has 100mph(V2=V1+V21=50+50=100)